A body is projected downwards at angle 30 to horijontal and velocity 9.8 from top oftower 29.4 m high how long it take befores strike ground

Grade:11

1 Answers

Bejoi Mathew
7 years ago
initial vertical velocity= usin30 = 4.9m/s
vertical distance traversed= 29.4m
So,
h=ut + 1/2gt2
29.4=4.9*t +4.9t2
on solving we get
t+ t – 6 = 0 (on solving quadratic equation)
t=3 seconds.
THANK YOU
(please note I have taken 9.8m/s2 as accleration due to gravity for simplicity)